OpenStudy (anonymous):

My brain is spacing.. Adam has 4 and 1/3 gallons of apple juice and 3 and 5/9 gallons of orange juice. How many gallons of juice does Adam have altogether in his refrigerator?

OpenStudy (anonymous):

71/9 simplify

OpenStudy (anonymous):

dont be mad at me sis

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

brainboozled please forgive me!

OpenStudy (anonymous):

during our break ill play roblox!

OpenStudy (salazarblack):

can anyone help me after this is done? http://openstudy.com/updates/53b848ebe4b09f140ca63c4f i have an exam tomorrow :'(

OpenStudy (anonymous):

pleeeeezzzeeee

OpenStudy (anonymous):

.-. @brainbuzzed5646 I'm going to do math now. But I will \[GLADLY\] play roblox at break with you.

OpenStudy (anonymous):

4 1/3 = 13/3 3 5/9 = 32/9 find a common denominator of 9 13/3 * 3 = 39/9 add em up then simplify you should get something like 7 1/9

OpenStudy (anonymous):

but that's an estimate fraction use your calculator

OpenStudy (anonymous):

kk so u \[FORGIVE\] me?

OpenStudy (anonymous):

\[PLEEZE\]

OpenStudy (anonymous):

\[Yes no maybe so\]

OpenStudy (anonymous):

\[now, GOODBYE.\]